OpenStudy (anonymous):

Hi, I'm just checking over an answer: 13= 5+3b-13 b= 7, correct?

jimthompson5910 (jim_thompson5910):

you are absolutely correct

jimthompson5910 (jim_thompson5910):

Check: 13= 5+3b-13 13= 5+3(7)-13 13 = 5 + 21 - 13 13= 26-13 13 = 13 which is true, so the solution is verfied

OpenStudy (anonymous):

correct!

OpenStudy (anonymous):

Thanks so much!